OpenStudy (anonymous):

Can someone please help me? I'm so confused. Use the quadratic formula to find the roots of the equation. Round to tenths if necessary. x2 - 6x + 2 = 0 {0.4, 5.6} {0.5, 4.8} {1.4, 2.4} no real number solution I keep doing the work but can't find the answer here is the work i did:

OpenStudy (anonymous):

I will wait to see what you post and tell you where, and if you went wrong. I will be lurking to see when you post your answer.

OpenStudy (anonymous):

here i did somethings. i hope u understand

OpenStudy (anonymous):

\[-1\pm \sqrt{1^{2}-4(1)(10)}^{}\] \[-1\pm \sqrt{1-39}\] \[-1\pm \sqrt{\frac{ 40}{ 2}}\] \[-1\pm \frac{ 40 }{ 2 }\] then i have two equations which are \[\frac{ -1+40 }{ 2 }\] and \[\frac{ 1+40 }{ 2 }\] they turn into \[\frac{ 39 }{ 2 }\] and \[\frac{ 1+40 }{ 2 }\] which is 19.5 and 20.5 which is apparently not right =/

OpenStudy (anonymous):

i know nit says 1-39 but i know that if i do put 1-39 taht it will be a negative number and i cant get the square root of a negative number =/

OpenStudy (anonymous):

ok, I have no idea where those numbers are coming from since your a b and c are not what you posted in the handwritten post. it reads \[x ^{2}-6x+2=0\]

OpenStudy (anonymous):

Where did you get your values from?

OpenStudy (anonymous):

holy crap your right!

OpenStudy (phi):

you are doing it correctly. the last step is round to the nearest tenth

OpenStudy (phi):

*in your hand-written notes*

OpenStudy (anonymous):

@phi im confused 19.5 and 20.5 are the answers? tahts not an option..

OpenStudy (anonymous):

can you write out teh steps and how you get the answer because im stumped...

OpenStudy (anonymous):

Those numbers you have in your typed response do not correlate to anything you have written emma,, so they must be from a different problem.

OpenStudy (phi):

oops, that was yunus post.

OpenStudy (anonymous):

haha ok lets start over....

OpenStudy (anonymous):

so the question was Use the quadratic formula to find the roots of the equation. Round to tenths if necessary. x2 - 6x + 2 = 0 {0.4, 5.6} {0.5, 4.8} {1.4, 2.4} no real number solution

OpenStudy (anonymous):

Ah...lol

OpenStudy (phi):

you need to pattern match a x^2 + b x + c =0 with your problem what is a? what is b? what is c?

OpenStudy (anonymous):

I was looking at two different posts!

OpenStudy (anonymous):

I think I been looking at numbers too long today.

OpenStudy (anonymous):

ok peoples i gtg to karate ill just have to figure this out some otehr timen than ks though fo rthe help :)

OpenStudy (anonymous):

the correct answer is 0,4 -5,6 just take a look what i did above.

OpenStudy (anonymous):

As i see you do so many wrong. it is almost all wrong what you did.